# thanks: This study was supported by the Russian Science Foundation (Project No. 23-62-10026) [RSF]

### v2

@@ Abstract
We present beam-dynamics simulations for an S-band RF photoinjector test bench under development at JINR, aimed at producing high-quality electron beams and enabling future generation of relativistic vortex electrons. Simulations of the 1.5-cell photogun under the presently available RF gradient of $\sim 45$ MV/m demonstrate stable bunch formation at low charge ($Q=0.63$ pC), where space-charge forces are weak and the transverse emittance is dominated by RF-induced correlations. Optimization of the injection phase and cathode-region solenoid yields a robust emittance-stabilization regime with a final normalized emittance of $\sim 3.2\ \pi\cdot\mathrm{mm\cdot mrad}$. To assess prospects for accelerating structured electron beams, we additionally model the quantum evolution of single-electron Laguerre–Gaussian wave packets. The results show that multi-MeV acceleration suppresses free-space spreading and preserves the wave-packet’s initial OAM structure, indicating that the test bench provides suitable conditions for forthcoming experimental studies of relativistic vortex electrons.
preprint: APS/123-QED
I Introduction
@@
The generation of relativistic vortex electron beams – quantum states with a helical phase front and a quantized orbital angular momentum projection (OAM) – is attracting growing interest across accelerator, atomic, and nuclear physics, electron microscopy, quantum optics, and foundation of quantum mechanics [Knyazev18, Bliokh17, Karlovets21, Ivanov22]. To date, vortex electrons have been generated only in the sub-MeV domain, up to kinetic energies of $\sim 300$ keV of transmission electron microscopes [Uchida10, Verbeeck10, Blackburn14]. Extending such structured states into the multi-MeV regime relevant for photoinjectors and linacs remains an outstanding challenge, yet promises significant impact, including new tools for nuclear and hadronic structure studies and potential alternatives to spin-polarized beams [Ivanov22].
@@
It is important to distinguish the genuine vortex beams of single particles whose wave functions are characterized by a helical phase structure $\exp(i\ell\varphi)$ and a quantized OAM projection $\ell\hbar$ from beams with modified transverse emittance ratios – such as flat and round beams demonstrated at Fermilab [Danilov96, Brinkmann02]. These latter beams can posses a so-called extrinsic angular momentum, which is a property of the beam as a whole, so a quantum state of each particle in the beam is not defined, whereas the former quantum beams of single electrons posses the so-called intrinsic angular momentum, the property of each electron akin to spin [Bliokh17, Floettmann20, Karlovets21, Ivanov22].
@@
One feasible route towards high-energy vortex electrons is the transfer of OAM from twisted ultraviolet light to electrons during photoemission [Karlovets21, Pavlov24, Kazinski25]. Whereas the generation of pure vortex states with a definite angular momentum projection and its vanishing quantum uncertainty is challenging due the finite temperature, space-charge effects, and other factors [Floettmann20, Karlovets21, Kazinski25], the generation of a superposition of vortex states with an uncertainty less than the mean OAM value of the beam seems feasible when the OAM is very large, $\ell\gg\hbar$, and the current is low so the space-charge effects can be mitigated. Realizing this mechanism requires a photoinjector capable of producing stable, low-emittance beams with well-controlled transverse profiles and rapid acceleration, ensuring that the initial OAM-imprinted phase structure of the emitted electron is preserved during the early acceleration.. To this end, a dedicated S-band RF photoinjector test bench is being developed at the Laboratory of Nuclear Problems, Joint Institute for Nuclear Research (JINR), combining a 1.5-cell RF gun [Nikiforov18] and a high-power UV laser system [Gacheva14]. Along with relativistic vortex electrons, there are plans at the Institute of Modern Physics in China to put a twist on relativistic ions [An25].
@@
RF photoinjectors are widely used as high-brightness electron sources at facilities such as LCLS [Alley99], the European XFEL [Ferrario01], and SuperKEKB [Cahill14], owing to their ability to quickly suppress space-charge effects via high-gradient acceleration [Giribono23, Maltseva24]. At JINR, there exists a LINAC facility, which is a linear accelerator of electrons to the energy of 200 MeV and plans to come to 400 MeV in 2027 [Nozdrin25], it operates with a thermionic DC gun and a multi-stage RF bunching system [Nozdrin21], producing $\sim 1$22 ps bunches. Replacing it with an RF photoinjector enables direct generation of $\sim 10$ ps bunches with the significantly improved transverse beam quality and precise control over the charge and spot size – critical prerequisites for producing structured quantum states.
@@
In this work, we present beam-dynamics simulations of the bunch generation and acceleration in the JINR photoinjector using realistic field maps. We demonstrate the implementation of a Carlsten-type emittance-compensation scheme: although the emittance does not decrease by an order of magnitude, the solenoidal lens stabilizes the transverse emittance at a low value and suppresses RF-induced phase-space distortions, establishing a robust operating point for subsequent vortex-beam experiments.
@@
We additionally investigate the quantum evolution of the single-electron Laguerre–Gaussian (LG) wave packets with different OAM values, up to $\ell=64\hbar$, accelerated in the same realistic DC and RF fields. The single-electron regime is justified by the planned experimental capability of operating at ultralow bunch charges approaching the quasi–single-electron limit. Collective effects such as Coulomb repulsion and decoherence are not included in this first study; thus the results represent an upper bound on coherence and OAM-structure preservation during acceleration.
II Test Bench Design
@@
The RF photoinjector test bench under construction at JINR is based on an S-band 1.5-cell RF photogun and a high-power UV drive laser. The overall layout (Fig. 1) includes the RF gun, the laser beamline, two solenoids for transverse beam control, and a 1.5-m vacuum transport section operating at $\sim 10^{-9}\ \mathrm{Torr}$.
Figure 1: Schematic layout of the RF photoinjector test bench. Key components include the S-band RF photogun, the UV drive laser, the solenoid lenses, the vacuum laser beamline, and the laser-injection mirror.
@@
The 1.5-cell S-band RF gun operates in the $\pi$-mode at 2856 MHz [Samofalova21]. Its internal geometry and photocathode placement (Fig. 2) are optimized for rapid extraction and emittance-compensating beam formation. The UV laser enters through the coaxial channel, which simultaneously serves as the RF coupling element and the electron-beam exit port. The on-axis accelerating field (Fig. 3) was obtained from CST Microwave Studio eigenmode simulations [CST].
Figure 2: Internal geometry of the S-band radiofrequency photogun. The design consists of 1.5 accelerating cells operating in the $\pi$-mode at 2856 MHz. The central aperture serves both for laser injection and electron beam extraction.
@@
Although the gun is designed for up to 6MW6\ $\mathrm${MW} of input RF power, the present test-bench operation is limited to about 3MW3\ $\mathrm${MW} because the cavity is still undergoing RF conditioning. Thus the $\approx 45\ \mathrm{MV/m}$ accelerating gradient used in the simulations corresponds to the experimentally achievable field at this commissioning stage.
@@
The emittance-compensation solenoid prototype (Fig. 4) was manufactured at JINR. It has an inner radius of 100 mm, outer radius of 140 mm, length of 100 mm, and contains 216 copper turns in six layers.
@@
The axial field profile, measured with 5-mm resolution at 20 A, is shown in Fig. 5. These measured values are used directly in the beam-dynamics simulations.
@@
III Beam Dynamics Simulation
All beam dynamics simulations were carried out with ASTRA [ASTRA], which couples a 3D particle-in-cell space-charge solver (in cylindrical symmetry) to high-order Runge–Kutta tracking through RF-cavity and magnetic fields. The key inputs are the bunch charge (from photocathode quantum efficiency and laser energy), mean initial energy and its spread, initial normalized emittance and spatial/momentum distributions, and the RF phase at photoemission.
@@ III Beam Dynamics Simulation
The photoemission model follows the Dowell–Schmerge approach [Dowell09], assuming Fermi–Dirac statistics and Gaussian spatial distributions. For an effective copper work function of $\phi_{\rm eff}=4.3\ \mathrm{eV}$ [Karkare15], the resulting mean kinetic energy and energy spread of emitted electrons lie in the sub-eV range, consistent with standard metallic photocathodes. With the UV laser spot size set to $\sigma_{x,y}=2\ \mathrm{mm}$ (equal to the initial RMS beam size), the corresponding intrinsic normalized emittance is
@@
A simulated RF-phase scan was performed to determine the optimal injection phase. Figure 6(a) shows the mean beam energy as a function of the initial RF phase $\varphi_{0}$. Maximum acceleration is achieved near $\varphi_{0}\approx 180^{\circ}$, with a phase tolerance of about $\pm 1^{\circ}$.
The phase scan of the normalized transverse emittance (Fig. 6(b)) shows a pronounced overall increase with injection phase in the explored interval. This behaviour is primarily determined by the intrinsic field structure of the 1.5-cell S-band RF gun operating in the $\pi$-mode: the axial electric field has a node close to the cathode plane, so the extraction field falls rapidly as the phase approaches the zero crossing. At the present low bunch charge (0.63pC0.63\ $\mathrm${pC}) the emittance is thus governed mainly by RF-induced transverse correlations imposed by the cavity geometry and mode structure, rather than by space-charge forces. This field configuration explains the observed monotonic phase dependence of the emittance. The operational working point of the test bench, chosen to maximize the energy gain under the available RF gradient, lies at $\varphi_{0}\approx 185^{\circ}$$190^{\circ}$. In this region the field on the cathode is not maximal, which inevitably increases the extraction-induced emittance, but it provides the highest attainable beam energy at the current stage of commissioning. The results of this scan define the initial parameters used in the subsequent beam-dynamics simulations (Table 1).
Figure 6: Simulated RF-phase scan results: (a) mean beam energy and (b) final normalized transverse emittance versus injection phase $\varphi_{0}$. Maximum acceleration occurs near $\varphi_{0}\approx 180^{\circ}$, while the operational working region $\varphi_{0}\approx 185^{\circ}$$190^{\circ}$ reflects the compromise between emittance and achievable energy.
@@
Normalized emittance 1.07  $\pi\cdot$ $m$ $\cdot$ mrad
@@
Optimal RF phase $185^{\circ}$
@@
III.1 Beam Dynamics Without Emittance Compensation
@@
Fig.7 shows the simulated mean kinetic energy of the bunch along the RF photogun beamline. A two‐peak structure arises from the axial field in the S‐band RF photogun, with rapid acceleration to $\overline{W_{\rm kin}}=1.9$ MeV at the gun exit, followed by a plateau. High‐gradient injection quickly suppresses space‐charge driven emittance growth once $\gamma\gg 1$, although early nonlinear distortions remain significant. Longitudinally, the bunch length $\sigma_{z}$ (Fig.8) increases rapidly during acceleration due to the velocity–energy correlation, as faster electrons at the head outpace slower ones at the tail; beyond the cavity, the expansion continues more gradually under residual space‐charge forces.
@@
The simulated energy evolution is defined by the presently available RF power. Whereas the gun can in principle deliver up to 5MeV5\ $\mathrm${MeV} electrons at its design RF-power limit of 6MW6\ $\mathrm${MW}, only 3MW3\ $\mathrm${MW} are available in the current hardware configuration. Consequently, the accelerating gradient is reduced to $\approx 45\ \mathrm{MV/m}$, resulting in a final kinetic energy of $\approx 1.9\ \mathrm{MeV}$ in Fig. 7. The reduced gradient also increases the relative importance of space-charge forces during the first few millimeters after the emission.
@@
Transversely, the beam expands symmetrically in $x$ and $y$ (Fig. 9(a)) from 2 mm to nearly 25 mm in the absence of external focusing. At the present low bunch charge (0.63pC0.63\ $\mathrm${pC}) and relatively large emission area, the space‐charge forces are moderate and contribute mainly to the divergence of the transverse envelope. The evolution of the normalized emittance, however, is dominated by correlated distortions from the transverse RF fields of the 1.5-cell $\pi$-mode gun, rather than by space charge. As shown in Fig. 9(b), the emittance remains near $1.5\ \pi\cdot\mathrm{mm}\cdot\mathrm{mrad}$ at the cavity exit but exhibits early oscillations characteristic of RF-induced correlations. These oscillations become particularly relevant when a solenoidal field is introduced, since the RF–solenoid coupling modifies their cancellation. A slight xxyy asymmetry likely originates from numerical sampling noise. Without external focusing, the beam size diverges rapidly and RF-induced emittance growth cannot be mitigated, motivating the use of a compensation solenoid in the following section.
@@
Figure 9: Simulated beam dynamics without emittance compensation. (a): Simulated RMS transverse beam sizes $\sigma_{x}$ and $\sigma_{y}$ along the beamline, showing symmetric growth from 2mm2\ $\mathrm${mm} to nearly 25mm25\ $\mathrm${mm}. (b): Simulated evolution of the normalized transverse emittance along the beamline. After initial RF-induced oscillations inside the 1.5-cell gun, the emittance remains approximately constant at the cavity exit and subsequently increases due to nonlinear RF-field–induced transverse phase-space distortions.
@@
Because the space-charge effects are weak in this regime, they can be neglected in the subsequent analysis. This is particularly important for the quantum-mechanical considerations in Sec. IV, since the small Coulomb forces do not significantly perturb the phase structure or the orbital-angular-momentum content of the evolving electron wave packet.
III.2 Emittance Compensation Using a Solenoidal Lens

### v3

# Classical and quantum beam dynamics simulation of the RF photoinjector test bench

@@ Abstract
We present beam-dynamics simulations for an S-band RF photoinjector test bench under development at the Joint Institute for Nuclear Research, aimed at producing high-quality electron beams and enabling future generation of relativistic vortex electrons with quantized orbital angular momentum (OAM). Simulations of the 1.5-cell photogun under the presently available RF gradient of 45 MV/m demonstrate stable bunch formation at low charge ($Q=0.63$ pC), where space-charge forces are weak and the transverse emittance is dominated by RF-induced correlations. Optimization of the injection phase and cathode-region solenoid yields a robust emittance compensated regime with a final normalized emittance of $3.2\ \pi\cdot\mathrm{mm\cdot mrad}$. To assess prospects for accelerating vortex electron beams, we additionally model the quantum evolution of single-electron Laguerre–Gaussian wave packets. The results show that multi-MeV acceleration suppresses free-space spreading of the electron packet and preserves the packet’s initial OAM structure, indicating that the test bench provides suitable conditions for forthcoming experimental studies of relativistic vortex electrons.
1 Introduction
@@
The generation of relativistic vortex electron beams – quantum states with a helical phase front and a quantized orbital angular momentum projection (OAM) – is attracting growing interest across accelerator, atomic, and nuclear physics, electron microscopy, quantum optics, and foundation of quantum mechanics [Knyazev18, Bliokh17, Karlovets21, Ivanov22]. To date, vortex electrons have been generated only in the sub-MeV range, up to kinetic energies of $\sim 300$ keV of transmission electron microscopes [Uchida10, Verbeeck10, Blackburn14]. Extending such structured states into the multi-MeV regime relevant for photoinjectors and linacs remains an outstanding challenge, yet promises significant impact, including new tools for nuclear and hadronic structure studies and potential alternatives to spin-polarized beams [Ivanov22, SchattschneiderPRB12, Wu19, An25].
@@
It is important to distinguish the genuine vortex beams of single particles whose wave functions are characterized by a helical phase structure $\exp(i\ell\varphi)$ and a quantized OAM projection $\ell\hbar$ [Bliokh17] from beams with modified transverse emittance ratios – such as flat and round beams demonstrated at Fermilab [Danilov96, Brinkmann02, Sun04, Piot06]. These latter beams can posses a so-called extrinsic angular momentum, which is a property of the beam as a whole, so a quantum state of each particle in the beam is not defined, whereas the former quantum beams of single electrons posses the so-called intrinsic orbital angular momentum, the property of each electron, complementary to spin [Bliokh17, Karlovets21, Ivanov22, Floettmann20].
@@
One feasible route towards high-energy vortex electrons is the transfer of OAM from twisted ultraviolet light to electrons during photoemission [Karlovets21, Pavlov24, Kazinski25]. Whereas the generation of pure vortex states with a definite angular momentum projection and its vanishing quantum uncertainty is challenging due the finite temperature, space-charge effects, and other factors [Karlovets21, Floettmann20, Kazinski25], the generation of a superposition of vortex states with an uncertainty less than the mean OAM value of the beam seems feasible when the OAM is very high, in our case up to $\ell=64\hbar$, and the bunch charge is low so the space-charge effects can be mitigated.
Realizing this mechanism requires a photoinjector capable of producing low-emittance beams with well-controlled transverse profiles and rapid acceleration, ensuring that the initial OAM-imprinted phase structure of the emitted electron is preserved during the early acceleration. To this end, a dedicated S-band RF photoinjector test bench is being developed at the Laboratory of Nuclear Problems, Joint Institute for Nuclear Research (JINR), combining a 1.5-cell RF gun [Nikiforov18] and a high-power UV laser system [Gacheva14]. Along with relativistic vortex electrons, there are plans at the Institute of Modern Physics in China to put a twist on relativistic ions [An25].
@@
RF photoinjectors are widely used as high-brightness electron sources at facilities such as LCLS [Alley99], the European XFEL [Ferrario01], and SuperKEKB [Cahill14], owing to their ability to quickly suppress space-charge effects via high-gradient acceleration [Giribono23, Maltseva24]. At JINR, there exists a LINAC facility, which is a linear accelerator of electrons to the energy of 200 MeV and plans to come to 400 MeV in 2027 [Nozdrin25], it operates with a thermionic DC gun and a multi-stage RF bunching system [Nozdrin21], producing $\sim 1$22 ps bunches. Replacing it with an RF photoinjector enables direct generation of 10 ps bunches in single-bunch operation, with the significantly improved transverse beam quality and precise control over the charge and transverse profile – critical prerequisites for producing structured quantum states.
@@
In this work, we present beam-dynamics simulations of the bunch generation and acceleration in the JINR photoinjector using realistic field maps. We demonstrate the implementation of a Carlsten-type emittance-compensation scheme. Although the emittance does not decrease by an order of magnitude, the solenoidal lens stabilizes the transverse emittance at a low value and suppresses RF-induced phase-space distortions, establishing a robust operating point for subsequent vortex-beam experiments.
We also investigate the quantum evolution of the single-electron Laguerre–Gaussian (LG) wave packets with different OAM values, up to $\ell=64\hbar$, accelerated in the same realistic DC and RF fields. We find that quantum transverse spreading in both accelerating field configurations is strongly suppressed compared with free-space drift, due to the rapid increase of the longitudinal momentum. This suppression enables realistic estimates of the transverse coherence length of electron wave packets in the photogun under experimentally achievable conditions.
@@
The single-electron regime is justified by the planned experimental capability of operating at ultralow bunch charges approaching the quasi–single-electron limit. Collective effects such as Coulomb repulsion and decoherence are not included in this first study; thus the results represent an upper bound on coherence and OAM-structure preservation during acceleration.
2 Test bench design
@@
The RF photoinjector test bench under construction at JINR is based on an S-band 1.5-cell RF photogun and a high-power UV drive laser. The assembled test bench (see Fig. 1) comprises the RF gun, the UV laser transport line, a steering magnet, two solenoids, and a 1.5-m vacuum beamline at $\sim 10^{-9}\ \mathrm{Torr}$.
Figure 1: Assembled RF photoinjector test bench during commissioning. The main components are indicated: RF gun, emittance-compensation solenoid, focusing solenoid, steering magnet, RF waveguide, UV laser transport line, vacuum mirror, and beamline.
@@
The 1.5-cell S-band RF gun operates in the $\pi$-mode at 2856 MHz [Samofalova21]. Its internal geometry and photocathode placement (see Fig. 3) are optimized for rapid extraction and emittance-compensating beam formation. The UV laser enters through the coaxial channel, which simultaneously serves as the RF coupling element and the electron-beam exit port. The on-axis accelerating field (see Fig. 3) was obtained from CST Microwave Studio eigenmode simulations [CST].
Figure 2: Internal geometry of the S-band RF photogun. The design consists of 1.5 accelerating cells operating in the $\pi$-mode at 2856 MHz. The central aperture serves both for laser injection and electron beam extraction.
@@
Although the gun is designed for up to 6MW6\ $\mathrm${MW} of input RF power, the present test-bench operation is limited to about 3MW3\ $\mathrm${MW} because the cavity is still undergoing RF aging. Thus the 45MV/m45\ $\mathrm${MV/m} accelerating gradient used in the simulations corresponds to the experimentally achievable field at this commissioning stage.
@@
The emittance-compensation solenoid prototype (see Fig. 5) was manufactured at JINR. It has an inner radius of 100 mm, outer radius of 140 mm, length of 100 mm, and contains 216 copper turns in six layers. The axial field profile, measured with 5-mm resolution at 20 A, is shown in Fig. 5. These measured values are used directly in the beam-dynamics simulations.
@@
3 Beam dynamics simulation
3.1 Definition of initial parameters
Beam dynamics simulations were carried out with $\mathrm{ASTRA}$ [ASTRA], which couples a 3D particle-in-cell space-charge solver (in cylindrical symmetry) to high-order Runge–Kutta tracking through RF-cavity and magnetic fields. The key inputs are the bunch charge (from photocathode quantum efficiency and laser energy), mean initial energy and its spread, initial normalized emittance and spatial/momentum distributions, and the RF phase at photoemission.
@@ 3.1 Definition of initial parameters
The photoemission model follows the Dowell–Schmerge approach [Dowell09], assuming Fermi–Dirac statistics and Gaussian spatial distributions. For an effective copper work function of $\phi_{\rm eff}=4.3\ \mathrm{eV}$ [Karkare15], the resulting mean kinetic energy and energy spread of emitted electrons lie in the sub-eV range, consistent with standard metallic photocathodes. With the UV laser rms beam size set to $\sigma_{x,y}=2\ \mathrm{mm}$ (equal to the initial RMS beam size), the corresponding intrinsic normalized emittance is
@@
Figure 6: Simulated dependence of the mean beam energy on the injection RF phase $\varphi_{0}$. Maximum acceleration is obtained near $\varphi_{0}\approx 180^{\circ}$, consistent with $\pi$-mode operation of the RF gun.
Figure 7: Simulated dependence of the final normalized transverse emittance on the injection RF phase $\varphi_{0}$. The operational working region $\varphi_{0}\approx 180^{\circ}$.
A simulated RF-phase scan was performed to determine the optimal injection phase. Figure 7 shows the mean beam energy as a function of the initial RF phase $\varphi_{0}$. Maximum acceleration is achieved near $\varphi_{0}\approx 180^{\circ}$, with a phase tolerance of about $\pm 1^{\circ}$.
The phase scan of the normalized transverse emittance (see Fig. 7) shows a pronounced overall increase with injection phase in the explored interval. This behaviour is primarily determined by the intrinsic field structure of the 1.5-cell S-band RF gun operating in the $\pi$-mode: the axial electric field has a node close to the cathode plane, so the extraction field falls rapidly as the phase approaches the zero crossing.
For realistic simulations, apertures corresponding to the internal structure of the RF gun were added (see Fig. 3): two apertures with an inner diameter of d=21.8mmd=21.8\ $\mathrm${mm} and a length of 10 mm, simulating the resonator cell walls, as well as an aperture with an inner diameter of d=14mmd=14\ $\mathrm${mm} and a length of 190 mm, simulating the coaxial input for impedance matching and RF power injection into the resonator cells. Simulations without the external solenoidal field demonstrate that most of the particles from the generated beam are lost within the first 25 cm from the cathode inside the coaxial input.
At the present low bunch charge (0.63pC0.63\ $\mathrm${pC}), the emittance is thus governed mainly by RF-induced transverse correlations imposed by the cavity geometry and mode structure, rather than by space-charge forces. This field configuration explains the observed monotonic phase dependence of the emittance. The operational working point of the test bench, chosen to maximize the energy gain under the available RF gradient, lies at $\varphi_{0}\approx 180^{\circ}$. In this region the field on the cathode is not maximal, which inevitably increases the extraction-induced emittance, but it provides the highest attainable beam energy at the current stage of commissioning. The results of this scan define the initial parameters used in the subsequent beam-dynamics simulations (Table 1).
@@
Normalized emittance $1.07\,\pi\,\mathrm{mm\,mrad}$
@@
Optimal RF phase $180^{\circ}$
@@
3.2 Beam dynamics without emittance compensation
@@
Fig.9 shows the simulated mean kinetic energy of the bunch along the RF photogun beamline. A two-peak structure arises from the axial field in the S-band RF photogun, with rapid acceleration to $\overline{W_{\rm kin}}=1.9$ MeV at the gun exit, followed by a plateau. High-gradient injection quickly suppresses space-charge driven emittance growth once $\gamma\gg 1$, although early nonlinear distortions remain significant. Longitudinally, the bunch length $\sigma_{z}$ (see Fig.9) increases rapidly during acceleration due to the velocity–energy correlation, as faster electrons at the head outpace slower ones at the tail; beyond the cavity, the expansion continues more gradually under residual space-charge forces.
The simulated energy evolution is defined by the presently available RF power. Consequently, the accelerating gradient is reduced to 45MV/m45\ $\mathrm${MV/m}, resulting in a final kinetic energy of 1.9MeV1.9\ $\mathrm${MeV} in Fig. 9. The reduced gradient also increases the relative importance of space-charge forces during the first few millimeters after the emission.
@@
Transversely, the beam expands symmetrically in $x$ and $y$ (see Fig. 11) from 2 mm to nearly 5.5 mm in the absence of external focusing. At the present low bunch charge (0.63pC0.63\ $\mathrm${pC}) and relatively large emission area, the space-charge forces are moderate and contribute mainly to the divergence of the transverse envelope. The evolution of the normalized emittance, however, is dominated by correlated distortions from the transverse RF fields of the 1.5-cell $\pi$-mode gun, rather than by space charge.
As shown in Fig. 11, the emittance remains near $1.5\ \pi\cdot\mathrm{mm}\cdot\mathrm{mrad}$ at the cavity exit but exhibits early oscillations characteristic of RF-induced correlations. These oscillations become particularly relevant when a solenoidal field is introduced, since the RF–solenoid coupling modifies their cancellation. Without external focusing, the beam size diverges rapidly and RF-induced emittance growth cannot be mitigated, motivating the use of a compensation solenoid in the following section.
Figure 10: Simulated RMS transverse beam sizes $\sigma_{x}$ and $\sigma_{y}$ along the beamline, showing symmetric growth from 2mm2\ $\mathrm${mm} to nearly 5.5mm5.5\ $\mathrm${mm}.
@@
Figure 11: Simulated evolution of the normalized transverse emittance along the beamline. After initial RF-induced oscillations inside the 1.5-cell gun, the emittance remains approximately constant at the cavity exit and subsequently increases due to nonlinear RF-field–induced transverse phase-space distortions.
@@
Because the space-charge effects are weak in this regime, they can be neglected in the subsequent analysis. This is particularly important for the quantum-mechanical considerations in Sec. 4, since the small Coulomb forces do not significantly perturb the phase structure or the orbital-angular-momentum content of the evolving electron wave packet.
3.3 Emittance compensation using a solenoidal lens